OpenStudy (anonymous):

Why does a P series converge if P is between 1 and zero??? doesn't it diverge but slowly?

OpenStudy (anonymous):

(assuming from n=1 to infinity)

zepdrix (zepdrix):

Hmmmmm it doesn't converge if the power is between 0 and 1. If you take for example the harmonic series, 1/n, it diverges, and that is where p=1. So in order for a P series to converge it needs to converge faster than the harmonic series (must have a P greater than 1). Hopefully I'm remembering that correctly :) lol

OpenStudy (anonymous):

Yes my books says that it will diverge but this makes no sense to me why does it have to be faster than the harmonic series? if I start doing 1/n^(1/2) for example, denominator gets bigger and bigger. why doesn't that mean that it will eventually converge?

zepdrix (zepdrix):

\[\frac{ 1 }{ \sqrt n }>\frac{ 1 }{ n }\] \[\frac{ 1 }{ \sqrt2 }>\frac{ 1 }{ 2 }\] Hmm I think the denominators are getting smaller actually. Making the actually terms get larger, meaning it doesn't converge quickly enough.

OpenStudy (anonymous):

\[\sum_{n=1}^{infinity} \frac{ 1 }{ n^\frac{ 1 }{ 2 } }\]

OpenStudy (anonymous):

n=1 : 1 n=2 : .707 n=3: .577 n=4: .5

zepdrix (zepdrix):

I think that only happens for the first few terms though, we want to know what happens long term. 1/10 = .1 1/sqrt10 = .316

zepdrix (zepdrix):

Hmm yah that is strange though :)

OpenStudy (anonymous):

so you agree that it seems like it should converge?

OpenStudy (anonymous):

i know that the answer is no but it annoys me that I don't know why

zepdrix (zepdrix):

If you were adding the first 7 or 8 terms, it would be less than 1/n, yes. But if you go any higher than that, the terms become much much larger than 1/n

zepdrix (zepdrix):

Did you see the example i gave with n=10? :O

OpenStudy (anonymous):

so the point is not if it is getting smaller but if it is larger or smaller than the harmonic at the corresponding term?

zepdrix (zepdrix):

You want to eventually be adding 0 with each term, and the size of each term that will get you there fast enough can be determined by the harmonic series. When p=1, 1/n grows too quickly to converge, the terms don't get small enough fast enough. When p=1/2, 1/n grows even FASTER than 1/n. Each term is larger than 1/n, so the series converses EVEN SLOWER than 1/n. Meaning it doesn't converge in the long run. For some reason I was getting confused by your decimals when you put the first few terms out there :) I looked at them wrong. For some reason I was thinking that the 4th term caught up to the 4th term in 1/n but it didn't.

zepdrix (zepdrix):

The terms need to be decreasing FASTER than 1 + 1/2 + 1/3 + 1/4 + ... + 1/n. In the case of the sqareroot, they're decreasing SLOWER than those terms. Still confused? D':

zepdrix (zepdrix):

 See how each new term you add is too large?